OpenStudy (anonymous):

What will the last payment be to pay off a $1,250 computer plus 8% sales tax with monthly payments of $120?

OpenStudy (anonymous):

Let's figure out what was payed for the computer. \[{\rm Cost} = 1250 + 1250 \cdot 0.08 = 1250(1.08)\] The final payment will be the remainder of the following division. \[Cost \over 120\]